July 24, e2e LIGO 1 meeting

 

 

e2e simulation group activity summary

 

 

e2e modeling

 

    1) LSC modeling by Luca

 

    2) Curvature mismatch

        * validation and bug fix completed for FP

        * moving toward PRM

 

    3) signal generation (Kathy : SURF student)

        * detector orientation and correlation

        * data file which e2e will read

 

----------------------------------------------------------------

 

Next actions

 

    1) Mechanics model

        * obtain ground motion spectrums of LHO and LLO (day&night) (Robert)

        * combine 6x6 stack transfer function

        * compare with Engineering run data to validate the model (Stan, Luca)

        * can the different of the locking status of LLO between day and night

          simulated ?

 

    2) Simulation of today's noise curve

        * noisy LSC

        * measured noise of later going into RM

        * calculate the noise curve from the dark port output

 

    3) PRM mode study

        * compare data, FFT and e2e to find how many modes are necessary

 

    4) Inclusion of mirror imperfections

        * GaryLynn will provide data by the end of August

        * lower order mode decomposition matrix

        * cavity asymmetry

 

    5) release of new e2e package by mid August

 

----------------------------------------------------------------

 

Alfi : Graphical User Interface of e2e

 

    A design and development of a new version has started.

    One major goal is to make it more stable and less picky about OS/versions

 

----------------------------------------------------------------

 

Brief summaries by Luca, Biplab and Virginio

 

    1) LSC modeling : Luca

 

    2) PSL noise study : Biplab

 

    3) Mechanics modeling : Virginio (quadruple adv. LIGO suspension)

 

 

Luca Matone : Modeling of realistic noisy LSC

 

List of measurements done so far on the 2k reflection port:

 

            o Demodulator Board (I)

                        - Noise floor measured;

            o Whitening Board

                        - Noise floor measured for different gain settings;

                        - TF with whitening filters on/off;

            o Anti-Aliasing Board

                        - TF with AA filter enabled (Nergis 2000);

            o Anti-Imaging Board

                        (- theoretical TF from schematics);

            o Digital filters

                        - TF measured and traced the corresponding TF coefficients of

                            + 0.1:10;

                            + 1k:10,100;

                            + 15:150;

                            + 1:20;

                            + 13HzRG;

                            + 1k:10;

                            + 2kBW;

                        - traced the filter coefficients of

                            + inverse of Whitening board;

                            + inverse of De-Whitening Board : dewhite1, dewhite2,

                              dewhite3, dewhite4;

                            + stopband (SB and Cheby);

            o Anti-Image (output) and Whitening Board (input) (PENTEK 6102)

                        - TF w/o Whitening filters on, w/o AA;

                        - Noise floor;

            o De-Whitening Board

                        (- theoretical TF from schematics and from inv(dewhite));

            o Noise floor of H2:LSC-REFL_I when

                        - light on PD is blocked;

                        - whitening filter enabled/disabled;

 

Measurements still to be done:

 

            o Anti-Imaging Board TF and noise floor;

            o Stopband digital filters TF;

            o De-Whitening Board TF and noise floor;

 

 

Biplab Bhawal : Study of PSL FSS noise due to higher order modes:

 (with Brad Zamft and Rick Savage)

 

* Studied the effects of the presence of higher order modes in input

  beam on frequency sensor noises.

* The Reference cavity of FSS of PSL is a symmetric one and so all

  higher order modes in both carrier and sidebands get reflected from

  it and fall on RFPD and beat with each other to give rise to signal

  which may be falsely assumed to be due to frequency shift and so

  acts as a source of noise.

* Studied the effects of PD motion or shift (longitudinal and transverse)

  and also nonuniformity in its responses with some realistic parameters.

* Our estimate, for example, in case of a 3 beam-width PD whose response

  function looks like a tilted surface with 1% variation across it and

  which is moving transversely with 1e-8 m/rtHz motion, is about one order

  of magnitude lower than the LIGO design requirement ( 1e-2 Hz/rtHz at 1 kHz).

  In other studied cases this level is lower and, especially for longitudinal

  motion of PD, it's of the order of the shot-noise limit (4e-5 hz/rtHz)

  or below it.

 

 

 

Detector Experts

 

 

Robert Schofield

 

Integration of environmental effects into the e2e model.

 

Spectra for normal seismic, acoustic, etc. environments

at LHO and LLO have been assembled into a document available at:

 

http://www.ligo.caltech.edu/~dhs/Adv-LIGO/T010074-01-D.pdf

 

Links to some original time series are also included.

 

Measurements of transfer functions to the interferometer have begun.

 

Day and night different << surf

correlation

tilt <<

 

Peter King

 

            Data gathering for the LLO PSL will commence shortly.  Until late

last night the laser was not available and earlier the laser was not

available outside the hours of 9pm to 8am.  In any case, I hope to finish

the following measurements before the end of this week.

            i.  laser free-running intensity noise

            ii.  relative intensity noise (RIN) just before suspended modecleaner

            iii.  RIN transmitted by the MC

            iv.  free-running frequency noise

            v.  frequency noise as measured by the MC - provided we manage to

                        pump down this week

 

 

 

David Reitze

 

IO e2e - Most of the activity at Hanford has been centered on restoring

the 2K MC to a quiet operating state (leveling), upgrading MC electronics

(installation of bounce mode notch filters, reparation of other

components).  High power operation has been tested and work; new

OSEMS are ~ 300 times less senstive to to 1.06 um light. Efforts at

LLO are underway to realign and relock the MC.